OpenStudy (anonymous):

Helppp!!!!....Indicate in standard form the equation of the line passing through the given points. M(0, 6), N(6, 0)

OpenStudy (mertsj):

The slope is -1 y-0=-1(x-6) x+y=6

OpenStudy (anonymous):

That is how you do the slope. y2-y1 ----- x2-x1